OpenStudy (anonymous):

A closed container has 4.02 ⋅ 1023 atoms of a gas. Each atom of the gas weighs 1.67 ⋅ 10–24 grams. Which of the following shows and explains the approximate total mass, in grams, of all the atoms of the gas in the container? 6.71 grams, because (4.02 ⋅ 1.67) ⋅ (1023 ⋅ 10–24) = 6.7134 5.69 grams, because (4.02 + 1.67) ⋅ (1023 ⋅ 10–24) = 5.69 0.67 grams, because (4.02 • 1.67) ⋅ (1023 ⋅ 10–24) = 6.7134 • 10–1 0.57 grams, because (4.02 + 1.67) ⋅ (1023 ⋅ 10–24) = 5.69 ⋅ 10–1

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (michele_laino):

we have to compute this: \[\Large \left( {4.02 \cdot {{10}^{23}}} \right) \cdot \left( {1.67 \cdot {{10}^{ - 24}}} \right) = ...?\]

OpenStudy (anonymous):

am i adding

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (anonymous):

Its D

OpenStudy (anonymous):

because you have to 4.02 + 1.67

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (michele_laino):

you have to multiply, here are your steps: \[\large \left( {4.02 \cdot {{10}^{23}}} \right) \cdot \left( {1.67 \cdot {{10}^{ - 24}}} \right)= \left( {4.02 \cdot 1.67} \right) \cdot {10^{23 - 24}} = ...?\]

OpenStudy (anonymous):

omg

OpenStudy (anonymous):

do i have to do this

OpenStudy (michele_laino):

yes!

OpenStudy (anonymous):

Its actually C

OpenStudy (anonymous):

you have tp multiply 4.02 * 1.67

OpenStudy (michele_laino):

yes!

OpenStudy (anonymous):

6 actually

OpenStudy (michele_laino):

4.02*1.67=6.7134

OpenStudy (anonymous):

6.7134 * 10-1

OpenStudy (michele_laino):

that's right!

OpenStudy (michele_laino):

which can be rounded off to 0.67 grams